# THÖR-MAGNI Act: Actions for Human Motion Modeling in Robot-Shared Industrial Spaces

### v1

@@ I Introduction
The recent THÖR-MAGNI dataset aims to address this gap by providing a large-scale indoor motion capture recording of human navigation and robot interaction [12]. It encompasses five distinct scenarios that simulate typical activities in industrial environments, such as the transportation of various objects, interaction with robots, and goal-oriented navigation alone and in groups. THÖR-MAGNI includes 3.5 hours of motion from 40 participants and 8.3 hours of egocentric videos from 16 participants over five recording days. The activity labels in THÖR-MAGNI are, however, limited to the invariable role of each participant (referred to as agent class), representing a complex activity assigned to the participant for the duration of the experiment. These roles have been shown to improve trajectory prediction [13] and describe gaze behavior [14], but fine-grained action labels would be required to model the separate sub-tasks in each activity.
@@
To demonstrate the utility of THÖR-MAGNI Act, we present two transformer-based prediction frameworks: (1) role- and action-conditioned trajectory prediction, extending [13], and (2) joint prediction of future trajectory and corresponding actions via multi-task learning [15]. Our results show how incorporating action labels can improve the performance of these predictive models. THÖR-MAGNI Act and the corresponding scripts are stored in a publicly accessible repository [^1] https://github.com/tmralmeida/thor-magni-actions. Documentation on how to use and visualize the dataset can be found in the same repository.
@@ II-A Experimental Design
The THÖR-MAGNI data acquisition, conducted in a laboratory, simulates industrial logistics settings to explore diverse human-human and human-robot interactions [12]. The dataset comprises five scenarios, distinguished by the spatial layout, the mode of the robot operation, and specific tasks assigned to the participants. These tasks or agent classes represent the high-level activities assigned to the person for the duration of a 4-minute recording session, such as: Carrier–Box, Carrier–Bucket, Carrier–Large Object, Visitors–Alone, Visitors–Group, and Visitors–Alone HRI. In each session, one to three participants wear eye-tracking glasses that capture egocentric video data. We provided standardized instructions to ensure natural behavior, informing participants that the experiment aimed to evaluate the robot’s perception of human actions [12]. In particular, participants were asked to carry stacks of boxes and buckets between the source and target locations (Carrier–Box and Bucket). Some participants moved a large poster stand (Large Object) in pairs of two. Visitors navigated freely between goal points, drawing a random card to determine their next destination. Visitor–Alone HRI involved both passive interactions (passing the robot in close proximity) and active interactions (joint navigation to goal points) with a mobile robot. The Carrier–Storage Bin HRI collaboratively with the robot transported storage bins between random goal points.
@@ II-B Action Annotations
In this work, we define a set $\mathcal{A}$ of 14 unique action labels: Walk, DrawCard, ObserveCardDraw (i.e., observing another person drawing a card at a goal point), WalkLO (i.e., moving a larger object), PickBucket, WalkBucket, DeliverBucket, PickBox, WalkBox, DeliverBox, PickStorageBin, WalkStorageBin, DeliverStorageBin, and HRI, based on the existing agent classes in THÖR-MAGNI. Each agent class is associated with specific actions, while some actions are shared across different agent classes (see Fig. 2 for an overview). An agent class is constant in all trajectories of a particular agent, whereas an action class may change at every time step. Consequently, this data extension provides finer labeling of internal factors (e.g., goal-driven actions) that can influence human motion. In particular, the action classes Walk and DrawCard are shared across multiple agent classes, indicating that trajectories involving these actions are likely to have similar characteristics, even when performed by agents of different agent classes.
@@ II-C Dataset Statistics
Figure 3: Top: 2D acceleration (mean $\pm$ one standard deviation), where values near zero indicate constant velocity. Middle: 2D velocity (mean $\pm$ one standard deviation), where values near zero correspond to static actions. Bottom: navigation distance (mean $\pm$ one standard deviation), where values near zero indicate static actions, and higher values reflect walking actions.
@@ III-B Multi-Task Learning for Trajectory and Action Prediction
The goal is to predict the future of a tracklet and the corresponding actions, $\psi_{\mathrm{MTL}}\colon(\mathbf{S}_{k},C_{k})\mapsto(\mathbf{Y}_{\mathbf{S}_{k}},\mathbf{a}_{\mathbf{S}_{k}})$, where $\mathbf{Y}_{\mathbf{S}_{k}}$ is the future tracklet and $\mathbf{a}_{\mathbf{S}_{k}}$ the future sequence of actions corresponding to the observed tracklet $\mathbf{S}_{k}$. The training data for this task, $\{(\mathbf{S}_{k},C_{k},\mathbf{P}_{\mathbf{S}_{k}},\mathbf{a}_{\mathbf{S}_{k}})\}_{k}$, consists of quadruples of observed tracklets, ground truth agent class labels, ground truth future positions and sequence of actions.
@@ IV Experiments and Results
Quantitative Results. Tab. I shows the results for the action-conditioned trajectory prediction task (TP) with various cues settings. It demonstrates that actions (third row) are powerful cues for trajectory prediction, outperforming the baseline (first row), with further improvements when combined with agent class information (last row).
@@ V Conclusions
The interplay between complex activities, actions, and locomotion dynamics of people and other agents still needs to be explored, particularly in industrial scenarios. This research gap can be addressed with comprehensive datasets that capture the relationship between actions and motion. Our work introduces the THÖR-MAGNI Act dataset to align action labels with diverse human motion cues. These cues, including position data, head orientation, gaze, and semantic attributes, provide a rich description of human motion in industrial settings. We also developed efficient and accurate transformer-based models for two key applications enabled by THÖR-MAGNI Act: (1) action-conditioned trajectory prediction and (2) joint action and trajectory prediction. THÖR-MAGNI Act with the diverse annotation classes pave the way for future research in human motion modeling based on rich contextual cues.
Acknowledgment
This work was supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) and by the EU Horizon 2020 No. 101017274 (DARKO).

### v2

@@ I Introduction
The recent THÖR-MAGNI dataset aims to address this gap by providing a large-scale indoor motion capture recording of human navigation and robot interaction [12]. It encompasses five distinct scenarios that simulate typical activities in industrial environments, such as transportation of various objects, interaction with robots, and goal-oriented navigation alone and in groups. THÖR-MAGNI includes 3.5 hours of motion from 40 participants and 8.3 hours of egocentric videos from 16 participants over five recording days. The activity labels in THÖR-MAGNI are, however, limited to the invariable role of each participant (referred to as agent class), representing a complex activity assigned to the participant for the duration of the experiment. These roles have been shown to improve trajectory prediction [13] and describe gaze behavior [14], but fine-grained action labels would be required to model the separate sub-tasks and their durations in each activity.
@@
To demonstrate the utility of THÖR-MAGNI Act, we present two transformer-based trajectory prediction frameworks: (1) role- and action-conditioned trajectory prediction, extending [13], and (2) joint prediction of future trajectory and corresponding actions via multi-task learning [15]. Our results show how incorporating action labels can improve the performance of these predictive models. THÖR-MAGNI Act and the corresponding scripts are stored in a publicly accessible repository [^1] https://github.com/tmralmeida/thor-magni-actions. Documentation on how to use and visualize the dataset can be found in the same repository.
@@ II-A Experimental Design
The THÖR-MAGNI data acquisition, conducted in a laboratory, simulates industrial logistics settings to explore diverse human-human and human-robot interactions [12]. The dataset comprises five scenarios, distinguished by the spatial layout, the mode of the robot operation, and specific tasks assigned to the participants. These tasks or agent classes represent the high-level activities assigned to the person for the duration of a 4-minute recording session, such as: Carrier–Box, Carrier–Bucket, Carrier–Large Object, Visitors–Alone, Visitors–Group, and Visitors–Alone HRI. In each session, one to three participants wear eye-tracking glasses that capture egocentric video data. In [12], we provided standardized instructions to ensure natural behavior, informing participants that the experiment aimed to evaluate the robot’s perception of human actions. In particular, participants were asked to carry stacks of boxes and buckets between the source and target locations (Carrier–Box and Bucket). Some participants moved a large poster stand (Large Object) in pairs of two. Visitors navigated freely between goal points, drawing a random card to determine their next destination. Visitor–Alone HRI involved both passive interactions (passing the robot in close proximity) and active interactions (joint navigation to goal points) with a mobile robot. The Carrier–Storage Bin HRI collaboratively with the robot transported storage bins between random goal points.
@@ II-B Action Annotations
In this work, we define a set $\mathcal{A}$ of 14 unique action labels: Walk, DrawCard, observing another person drawing a card at a goal point (ObserveCardDraw), moving a larger object (WalkLO), PickBucket, WalkBucket, DeliverBucket, PickBox, WalkBox, DeliverBox, PickStorageBin, WalkStorageBin, DeliverStorageBin, and HRI, based on the existing agent classes in THÖR-MAGNI. Each agent class is associated with specific actions, while some actions are shared across different agent classes (see Fig. 2 for an overview). An agent class is constant in all trajectories of a particular agent, whereas an action class may change at every time step. Consequently, this data extension provides finer labeling of internal factors (e.g., goal-driven actions) that can influence human motion. In particular, the action classes Walk and DrawCard are shared across multiple agent classes, indicating that trajectories involving these actions are likely to have similar characteristics, even when performed by agents of different agent classes.
@@ II-C Dataset Statistics
Figure 3: Top: 2D acceleration (mean $\pm$ one standard deviation), where values near zero indicate constant velocity. Middle: 2D velocity (mean $\pm$ one standard deviation), where values near zero correspond to static actions. Bottom: navigation distance (mean $\pm$ one standard deviation), where values near zero indicate static actions and higher values reflect walking actions.
@@ III-B Multi-Task Learning for Trajectory and Action Prediction
The goal is to predict the future of a tracklet and the corresponding sequence of actions, $\psi_{\mathrm{MTL}}\colon(\mathbf{S}_{k},C_{k})\mapsto(\mathbf{Y}_{\mathbf{S}_{k}},\mathbf{a}_{\mathbf{S}_{k}})$, where $\mathbf{Y}_{\mathbf{S}_{k}}$ is the future tracklet and $\mathbf{a}_{\mathbf{S}_{k}}$ the future sequence of actions corresponding to the observed tracklet $\mathbf{S}_{k}$. The training data for this task, $\{(\mathbf{S}_{k},C_{k},\mathbf{P}_{\mathbf{S}_{k}},\mathbf{a}_{\mathbf{S}_{k}})\}_{k}$, consists of quadruples of observed tracklets, ground truth agent class labels, ground truth future positions and sequence of actions.
@@ IV Experiments and Results
Quantitative Results. Tab. I shows the results for the action-conditioned trajectory prediction task (TP) with various cues settings. It demonstrates that actions (third row) are powerful cues for trajectory prediction, outperforming the baseline (first row), with further improvements when combined with agent class information (last row). The additional parameters result from the action classes included in the input layer.
@@ V Conclusions
The interplay between complex activities, actions, and locomotion dynamics of people and other agents still needs to be explored, particularly in industrial scenarios. This research gap can be addressed with comprehensive datasets that capture the relationship between actions and motion. Our work introduces the THÖR-MAGNI Act dataset to align action labels with diverse human motion cues. These cues, including position data, head orientation, gaze, and semantic attributes, provide a rich description of human motion in industrial settings. We also developed efficient and accurate transformer-based models for two applications where THÖR-MAGNI Act can play an important role: (1) action-conditioned trajectory prediction and (2) joint action and trajectory prediction. THÖR-MAGNI Act with the diverse annotation classes pave the way for future research in human motion modeling based on rich contextual cues.